These are the news starting 01 Jan 2012
Click here for the news of the previous years.

31 Dec 12 Uploaded overlay of  total ozone column graphs from Uccle and meteoLCD  Our measurements seem a bit too low, but reflect well the common spikes and lows. A more detailed analysis will be made asap.
This is the last news for this year.

As a recall: MeteoLCD started in 1994 and became fully operationel in 1998, which is the base year from which we compute several of our trends.

17 Dec 12 Added DOBSON (total ozone) measurements made by Mike Zimmer during November. There was a problem with the graph making procedure (due to my working with different html editors), which is now corrected.
14 Dec 12 Maintenance work by T.S.E.
12 Dec 12 Uploaded nov12.dat file to data archive.
11 Dec 12 1. Download of November 2012 data
2. Replaced running UVB sensor (#5760) with model #1894 at 09:45 UTC, which is on return from Solar Light Co. where it has been recalibrated. The new sensitivity is down by 4% in respect to the previous calibration of the #1894, (which operated until 10 Sep. 2012 ) what means that the new readings will be slightly lower. As the systematic calibration error is between -6.2 and +4.7%, this will not have a big relevance..
14 Nov 12
There was a solar eclipse today viewable in large parts of Australia. Here a picture I took at 08:00 local time with a rudimentary equipment ( a looking glass!) of the eclipse, close to the maximum of about 50% as viewable from a location in Victoria (click to enlarge):
partial solar eclipse,
                Foster, Australia (VIC)
06 Nov 12
Ground ozone sensor is back from calibration and went online today.
01 Nov 12 Uploaded oct12.dat file into data archive; closed running all.dat file and uploaded as all_05Nov11to31Oct12.dat. New all.dat will start at midnight.
28 Oct
1. Maintenance work. As no CO2 span gas is available, our zero calibration could be done.
2. The heat stress WBGT sensor has been stopped.
3. The ground ozone sensor will probably be back during the next week. The icons and plots of the live data page have been returned to the original to show ozone and air temperature, just to be sure to correctly display everything when the ozone sensor is working again.
So do not be bewildered by the curious -560 ug/m3 ozone readings  when there is yet no sensor available..
13 Oct 12 Uploaded sep12.dat file into data archive
18 Sep 12 Recalibration of the AC31M (NOx) sensor. Cleaning of the terrace from the numerous bird droppings. (yes, that's how real science works:-))
14 Sep 12 Maintenance and calibration by T.S.E. The NOx sensor is again operational.
Attention: The ozone sensor has been removed and will be recalibrated at Brussels. The sensor will be unavailable for quite a long time, probably up to November!
10 Sep 12 1. Replaced UVB sensor # 1894 by #5760 at 07:30 UTC. The older sensor will be sent back to Solar Light Co. for recalibration. Model #5760 was last calibrated in Dec. 2001. A comparison done in May 2012 showed that the new UVB5760 readings should be multiplied by 0.896 to make them comparable to those of the model #1896 (see calibration graph)
2. Uploaded aug12.dat file into data archive
22 Aug 12 The radioactivity readings went bad  the 12th Aug 2012. probably due to a computer failure. Restarted today at 07:10 UTC; seems ok.
09:16: Well, the problem was more serious that anticipated. After some time, the computer hangs, and rebooting does not cure the problem. Replaced computer by a clone ( I had a recent image of the original HD). Ok now....
01 Aug 12 1. Uploaded jul12.dat file into data archive
2. There are two important problems:
- the NO/NO2 sensor is out order, and will not being repaired befor mid-September
- there is an intermittent problem with the temperature regulation (UVB-Tin) of the UVB sensor, starting the 21 Jul 2012 at 04:00.UVB readings (mMED/h) after that date should be accepted with caveats.
26 Jul
We now have again the usual ground ozone peak, a regular phenomenon when the sky is blue,UVB and temperatures high. Two comments about that:
1. I am pretty sure that the high rural concentrations (that are used to launch mandatory traffic slowing to 90 km/h on the autobahn's) are mostly caused by VOC's (terpenes, isoprenes...) naturally emitted by conifers.
Reducing traffic speed on  the Luxembourg highways (which are far away from the reference station in Vianden) will have no measurable effect.
2. There is a short interview with me at the journal of Radio RTL (link here, spoken language is Luxembourgish; jump to time stamp10:50)
3. See a poster session "10 years ozone measurements at meteoLCD" discussing these problems
4. I think that our actual ground ozone readings at Diekirch are too low by about 10 -20%. In September we will try  to bring the instrument to Brussels for a general recalibration
09 Jul 12 1. The NOx sensor (AC31M) definitively has gone bad: there is an ozone generator fault I am unable to correct. The sensor will be switched off, probably up to mid-September :-)))
2. The lightning movies seem to be made correctly again ...
08 Jul 12 There is again a problem with the NOx sensor, which stopped working properly during the 06th July 2012. We did not find a clear cause for this malfunction., which sometimes miraculously disappears after some fiddling and down/up powering...
07 Jul 12 There remained some serious problems with making the daily lightning movie. If ffmpeg was called by the "convert" command from ImageMagick, things went sometimes spectacular wrong, with pictures from foreign directories being incorporated... I changed the procedure by not using Imagemagick anymore and making the mpeg video using only ffmpeg. Tests done this morning (09:40) suggest that this works correctly now..
04 Jul 12 Uploaded June 2012 data into the data archive.
03 Jul 12 The LNI zero air device went bad (not enough pressure delivered). Raoul Tholl and myself worked for over two hours to find the problems: one very dirty water trap filter and a perforated tube which leaked air flow to the outside. Repaired at 10:00 UTC.
19 Jun
Meteo01 and lightning stuff running again. I installed a new system disk and had to change the codec for making the movies...A lot of work!.
18 Jun 12 This definitively is a BLACK MONDAY!
Detective work to find some errors:
1. the logger stopped communicating at 03:30 (obscure COM1 error: solved by rebooting...hmm!)
2. the NOx sensor now seems ok
3. the NOx icons were not drawn correctly (error in a an FTP command file)
4. some interruption from 08:15 to 09:15 because of imaging full hard disk of the Linux server making the graphs This really had to be done!

5. The main hard-disk of the meteo01 server (Linux) went bonk (probably after some power outages, despite UPS). This server handles most of the lightning related stuff, including the making of the movies. I moved the main page displaying the "radar-screen" to our server, and disabled the links to the movies (ok at 13:15 UTC). The repair of the dead server is quite a lot of work, thank you for not expecting wonders!

17 Jun 12 1. New paper by F. Massen and 5 co-authors on radioactivity measurements at meteoLCD. See here.
2. There is a serious problem with the NOx sensor. For the moment, readings should be discarded.
15 Jun 12 Calibration and maintenance work on the gas sensors done by Raoul Tholl. NOx sensor readings should be taken with suspicion!
10 Jun 12 Uploaded May 2012 data into the data archive.
20 May 12 UVB compare action stopped at 06:50. The UVA sensor is reinstalled, and the UVA graph will from now on show again the correct UVA irradiance in Wm-2. ( from 14 to today the legends were UVA but what was measured wrtr the mMED/h readings of the #5760 sensor).
Attention #1: the scale of the plots (today and previous days) will adjust automatically. If you want the UVA data, download the ASCII files.
Attention #2: To revert to the original configuration, I had to stop the datalogger, download all data, make the config. changes and restart. This has been done from 11:32 to 11:40, and the logger will resume at 12:00. The plots and data of channel 10 (UVA) will start with hopefully correct values at that moment!
A quick research shows that a multiplier of 0.8955 should be applied to the #5670 readings to make them comparable to those of instrument #1894.
16 May 12 Yann Logelin, the presenter of the regular weatherbulletin (METEO) on the national RTL TV station showed today a screen with the UVI measurements of meteoLCD, including all references and clearly mentioning the source of his information. This is really appreciated! See replay of emission here. and screenshot here.
14 May 12 1. Recalibration of gas sensors (R. Tholl and F. Massen)
2. The Davis VantagePro backup station had a loose windvane/anemometer arm. Fixed. Check of windvane shows that true North gave reading of 319° (instead of 360 or 0). I am not sure for haw long this offset has been active. Recalibrated and checked for correct reading of North direction.(09:00)
3. Replaced pump of zone sensor (15:00)

4. Important: We plan to send the UVB sensor (#1894, channel 9) back to Solar Light Co. for recalibration. When off, this sensor will be replaced by #5760 (same type). To make a prior intercomparison between the 2 UVB sensors, the UVA sensor at ch. 10 is temporarily replaced by the UVB sensor #5760. So the UVA plot shows the mMED/h of the sensor #5760 (discard the wrong text saying UVA). This comparison will run for about 4 days.

06 May 12 Uploaded Apr12 data to data archive
04 May 12 1. Maintenance by T.S.E
2. Restarted WBGT measurements at 15:30 UTC
3. Installed new NAS server for storing the images of the many computers (disk images used for cloning).
24 Apr
The Radioprotection Office installed a self-contained dual Geiger counter on the meteoLCD terrasse. The device is powered by a solar panel and transmits its readings by GSM phone. It will be used to make an comparison with our radioactivity measurements. If you use the webcam and choose UV_sensors you may see a vertical cylinder right of the UVB (the left-most) sensor. That's the upper part of the Geiger counter.
01 Apr 12 Uploaded Mar12 data to data archive
29 Mar 12 Recalibration and maintenance. Thorough cleansing of solar sensors and air sample inlet.
06 Mar 12 The Luxembourg Radio Station RTL had a interview over phone with German author Fritz Vahrenholt (Die kalte Sonne) and myself. Click here for the first file (in German, wma) and here for the second (in Luxembourgish, wma)
05 Mar 12 Maintenance and recalibration (together with Raoul Tholl)
04 Mar 12 1. uploaded Feb12 data to data archive
2. The Delta_T logger does not handle leap years; this has been manually corrected in the feb12.dat file. The logger has been halted, emptied, date reset and restarted. No data will be lost.
10 Feb 12 1. replaced ozone sensor pump by one we (technician J-P. Klein an myself) have refurbished (all 3 ball bearings replaced). Pump now runs very smoothly! This home-made repair made us an economy of over 1200 Euros!
2. Recalibration of AC31 NOx sensor.
05 Feb 12 1. Uploaded jan12.dat file into data archive.
2. It is a normal situation that there are no moist enthalpy data: moist enthalpy can be computed only for positive air temperatures, and we are in deep in a cold snap! It might seem inelegant that GNUPLOT does not create anything picturesque in that case, but I think you should bear this.
03 Feb 12 Maintenance by TSE. The Zero Air generator is still out of order, it was impossible to repair the compressor. The complete module compressor/motor must be replaced.
13 Jan 12 1. The bearing of the O3 sensor pump was nearing death... pump replaced by a refurbished in stock. Makes also some disquieting noises, hope the pump will last some time.
2. I made a quick and dirty repair of the zero air  LNI generator: as we have not yet a tool small enough to remove the main bearing (tool is ordered), I exchanged the motor against an old one, a leftover from a previous repair, keeping the original compressor head. Seems to be ok (16:30), so I am making the usual calibration of the CO2 sensor.

...well it did not last long: I just was able to finish the CO2 calibration, when the compressor stopped again.!

So definitive repair of the zero air generator will be made during the next visit of Envicontrol (26 Jan 2012).

09 Jan 12 Partial recalibration and maintenance with colleague Raoul Tholl. The LNI zero air generator has broken down (bad bearing in compressor motor). Repair will not be immediate. This should not have big consequences, as both CO2 and O3 sensors have their own zero air facility. There just should be not people standing near the CO2 sensor, as this instrument must constantly be flushed by CO2 free air (what a bad design by the constructor API Teledyne!)
08 Jan 12 TRENDS update nearly finished. NOx data remain to be updated.
07 Jan 12 Started work with updating TRENDS section to 2011 data. This is work in progress!
02 Jan 12 1. Uploaded dec11.dat file into data archive
2. Also uploaded = data file containing all months, including meta data
3. And made the traditional 2011_only.xls file with all the data of the year 2011 and the usual statistics.
01 Jan 12 The usual new year problem with the 7 days graphs.. as the logger file does not contain the year, Gnuplot gets confused the 1rst week of every new year. This will self correct the 7th January.

Remember that the ultrasonic anemometer is out of action (possibly  for a longer time). Take wind velocity from ch. 61 (labeled Airspeed) and add an offset of 0.38 m/s to make readings of this lower mounted cup anemometer comparable to the ultrasonic anemometer at ch. 4 (labeled Air_Velo) Wind direction should be read from ch.13 (labeled Windvane); ch. 15 (labeled WindDir) corresponds to the ultrasonic anemometer, and should also be discarded.
So in short, here are the channels from left to right in the data file that should be discarded:
channel Label Comment channel to use instead
4 Air_Velo ultrasonic anemometer out of action ch. 61
7 UVA old Delta_T sensor ch. 10
8 CO2 old MIR9000 sensor ch.11
15 WindDir ultrasonic anemometer out of action ch. 13
23 Solarrad old Delta_T solar sensor ch. 6

The main reason for keeping the data files unchanged is that a lot of users pick data from these files and rely on a stable format.




Back Home